I have to multiply (3+ 2i)(2-5i)7i how do I do this? I have no clue where or how to begin.
thanks in advance!

Answer by Alan3354(69443)

multiply (3+ 2i)(2-5i)7i
---------
Do the binomials:
(3+ 2i)(2-5i)
Do it the same as any binomial multiply.
(3+ 2i)(2-5i) = 3*(2-5i) + 2i*(2-5i)
= 6-15i + 4i-10i^2
i^2 = -1
= 6-15i + 4i+10
= 16 - 11i
---------
Multiply by 7i
7i*(16 - 11i) = 112i - 77i^2
i^2 is still = -1
= 112i + 77
= 77 + 112i